Main INDEX
Monthly INDEX
PREV
NEXT
User name gilman
Log entry time 00:12:07 on June 3,2000
Entry number 42908
keyword=evening shift summary
No beam , waiting for injector experts to come in, then Moller when beam back
Jack warns: EPICS problem, do NOT NOT NOT...(odd number of times)
reboot IOCs, things will stop working.
17:20 MCC says injector people's work done in 30 minutes
17:45 Jack calls and says we can reboot IOCs as we please!
20:53 serious beam starts coming into the hall, Moller gets started!
22:00 Moller ~ done when beam goes away
22:20 pulsed beam allows deflection test with Moller, start switchover back
22:42 start taking data: run 1695 carbon raster on
22:50 start run 1696: carbon raster off, and queue through dummy targets
for pointing tests...
23:30 5 uA on LH2 to check raster, run 1700
try taking data at 50 uA, S/N looks lousy on line